Find the electric field strength due to a point charge of 5 micro coloum at a distance of 80 cm from the charge

Grade:12

2 Answers

Arun
4 years ago
 
As we know
E = V * r
Hence
E = (1/4* pi* epsilon) * (q/r)
E = 9 * 10^9 *5 * 10^-6 /0.8
E = 45000/0.8
E = 56250 volt - m
Khimraj
4 years ago
 
Electric field due to a point charge is given by
E = Kq/r2
K = 9*109
q = 5*10-6C
r = 0.8m
SO 
E = 7.03*104 Vm.
Hope it clears.